State the formula for the phenotypic variance in a population, give a brief definition of each term?

Interpret the meaning of an H2 value (broad-sense heritability) that approaches 1.0

Interpret the meaning of the H2 value (broad-sense heritability)that approaches 0.0
Answer #1

Phenotypic variance (Vp) is the sum of variance caused by genetic (Vg) and environmental factors (Ve).

It is calculated as:

Vp = Vg+ Ve

The genetic variance may includes dominance variance (Vd), additive variance (Va), and epistatic variance (Vi).

Vg= Vd+ Va +Vi

This gives:

Vp =Vd+ Va +Vi + Ve

Genetic variance results in a variance in phenotypes because of different alleles, mutations, etc. Dominance variance is caused by alleles interacting at one specific locus. On the other hand, the epistatic variance is caused by differential interaction of alleles at different loci. Additive variance is the sum of the effects of two alleles interacting together to form a different phenotype.

Different environmental conditions (pH, temperature, altitude, sunlight, sea depth, etc.)  result in variations called environmental variation.

Heritability (H2) is the ratio of genetic variance and phenotypic variance.

H2= Vg/Vp

Heritability value ranges from 0 to 1.

An H2 value 1 or close to 1 indicates the presence of more genetic variations. The H2 value 0 indicates that genetic variation is not present in a population.